# Rapid Bayesian identification of sparse nonlinear dynamics from scarce and noisy data

### v1

@@ 1 Introduction
Another prevalent problem with existing SINDy-like algorithms is their ability to handle scarce and noisy data. While SINDy-based approaches are effective at identifying equations from synthetic data with little noise, real-world measurements are typically scarce and prone to significant measurement error. Large noise in the data presents several challenges to SINDy. First, measurement noise corrupts the precise recovery of time derivatives because conventional finite difference approximations can amplify noise by the order of the sample rate. This problem can be mitigated by using a weak formulation [13, 14]. However, this approach may also introduce bias to the data, as we will illustrate in §2.3. Second, excessive noise in the data can impede STLS’s ability to select the most appropriate model. STLS’ poor performance in noisy and scarce data is partly due to the fact that it does not account for the uncertainty in the data, even though this uncertainty can be a crucial factor to consider when choosing a model. Furthermore, the performance of STLS deteriorates when the number of data points is lower than a certain threshold. As we will demonstrate later in §4.1, by adopting a Bayesian viewpoint, one can better find the reason for such deterioration and the remedy against it, effectively addressing the constraints posed by scarce and noisy data.
@@
Therefore, the goal of this work is to demonstrate how a Bayesian perspective can encompasses all preceding techniques and is best suited for extending the current SINDy techniques to function in challenging data environments where data are noisy and scarce. As we shall demonstrate later in §4.1, by grounding SINDy in the Bayesian framework, we can also more rigorously justify how and why SINDy works well in the high data low noise limit.
@@ 1.1 Related work and contribution of this work
Sparse Bayesian learning (or regression) is a mature and well-established method in the statistical literature [19, 22, 23]. The dictionary learning technique typically employs a sparsity-inducing prior to promote sparsity. The current ‘gold standard’ sparsity-inducing priors are the spike-and-slab-typed priors, but earlier work has also investigated the use of hierarchical priors. For example, when one combines a Gaussian prior with variance $\alpha^{-1}$ with a maximum-entropy hyperprior on $\alpha$ (i.e. a scale uniform distribution), the result is a sparsity-inducing student-t true prior that promotes sparsity. A notable early study that utilises this hierarchy prior to promote sparsity is the RVM/SparseBayes package developed by Tippings & Faul [24]. The Relevance Vector Machine (RVM) is a type-II-maximum-likelihood algorithm that was initially created as a probabilistic alternative to the Support Vector Machine. It utilises kernels as the dictionary and aims to maximise evidence while also encouraging sparsity in dictionary learning. This method exploits the fact that the conditional hyperprior given the data can be approximated as a Dirac Delta, while the rest of the probability distributions can be approximated as Gaussians, which allows for analytical tractability. The resulting approach has been demonstrated to be both computationally efficient and robust, not only for classification tasks but also for addressing general sparse regression problems. Due to its Bayesian foundation, it is also resilient to noisy and scarce datasets. Therefore, it is especially appropriate for the objective of this study.
@@
We note that this work is not the first to apply sparse Bayesian regression in the context of equation discovery or system identification [25, 26, 27], nor the first to integrate sparse Bayesian regression with SINDy [28, 4]. For instance, Zhang & Lin [28] incorporated an additional thresholding procedure into the SparseBayes method and applied the technique to extract different ordinary differential equations (ODEs) and partial differential equations (PDEs) from numerical data. They conducted a comparative analysis of their threshold sparse Bayesian regression algorithm with STLS and LASSO, and demonstrated Bayesian regression’s superior performance in the presence of noisy data. Recent work by Niven et al. [27] has also attemtped to use Bayesian inference for model selection and uncertainty quantification in system identification. However, the synthetic dataset they considered only introduced uncertainty in the time derivatives and not in the state variables themselves, which can oversimplify the challenge of the system identification problem, as we shall demonstrate later in (9). Fuentes et al. [4] substituted STLS with SparseBayes directly as the sparse regression algorithm and employed the technique for system identification. In particular, apart from synthetic data, they have successfully applied this approach to two experimental benchmarks. Nevertheless, the dataset they used has a high sampling rate (i.e. a high data limit). When finite difference was applied, the noise was significantly amplified in their time derivative, resulting in unreasonably large uncertainty in their learning result despite successfully recovering the desired equations.
@@
While the above work uses variants of SparseBayes, which do not require expensive computation, they rely on Gaussian approximations of the likelihood and prior distribution in order to maintain the analytical tractability of the distributions. In contrast, subsequent works [29, 30, 31, 32] use spike-and-slab-typed priors, which, despite the more accurate recovery of the correct model, necessitates costly sampling techniques, such as MCMC [29, 31] or Approximated Bayesian Computation [30] to capture the distribution of the estimated parameter more accurately. The benefit of using these sampling-based techniques lies in their capacity to precisely capture non-Gaussian distributions in the parameter and their ability to integrate sparsity-promoting spike-and-slab-typed prior. However, they come at the expense of expensive computations. A cheaper alternative is to make use of ensembling technique [12] to estimate uncertainty in the model. Recent work [33] has demonstrated the convergence of the bootstrapping-based STLS algorithm to the equivalent Bayesian statistics. Nonetheless, these algorithms remain an order of magnitude more expensive than STLS and the Gaussian approximations used by SparseBayes.
@@
Given that our goal is to seek a fast and computationally inexpensive learning method, this work will follow the earlier work more closely [28, 4] and shall adopt strategies similar to the SparseBayes method. Despite its similarity with the earlier work, we note a few key differences. First, our work highlights the key strength of the Bayesian framework, namely, its superior performance in scarce and noisy data. This is particularly in contrast to [4], whose dataset is in the high data limit. Second, our work compares the performance of the weak formulation and finite difference in obtaining derivatives from scarce and noisy data and highlights the pros and cons of each method, which is in contrast to the finite difference used in earlier work [28, 4]. Lastly, we shall present a new sparse Bayesian regression algorithm, which, despite heavy inspiration from SparseBayes, gives better performance due to its ability to keep track of how noise variance propagates throughout the calculation.
@@ 1.2 Structure of the paper
The rest of the paper is structured as follows. In §2, we will provide a concise overview of the SINDy framework. The significance of noise will be emphasised, particularly in terms of its propagation through the collection of potential functions and the computation of temporal derivatives. Then, we will introduce our Bayesian approach, which takes into account the propagation of noise, and briefly compare our Bayesian algorithm with previous studies. We will also briefly highlight the difficulty of approximating time derivatives using weak formulations when the sampling rate is low, which further reinforces our case for a noise-robust algorithm. In §3, we will showcase the efficacy of the suggested framework by presenting examples such as the Van Der Pol system, the nonlinear oscillator, the Lorenz system, and predator-prey dynamics. The predator-prey dynamics are based on real-life data, highlighting the practicality of the proposed framework in real-world applications. Then, in §4, we will analyse the reasons why the Bayesian approach is more resilient in scarce and noisy data and highlight some limitations of the proposed algorithm. In particular, in §4.1, we will explain why SINDy works in the high data limit from a Bayesian perspective, but not as well as its Bayesian counterpart in the low data limit. Finally, we will discuss how the improved data efficiency and the probability framework create new potential innovations in active learning.
@@ 2.1 SINDy and tracing noise propagation
Suppose we have a system of interest with state variables $\hat{\mathbf{x}}=[\hat{x}_{1}(t),\;\hat{x}_{2}(t)\;...\;\hat{x}_{D}(t)]\in\mathbb{R}^{1\times D}$ with dimension $D$, governed by the dynamical equation
@@
with some unknown function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. The system was observed $\tilde{N}$ times at $t=t_{1},\;t_{2},\;...t_{\tilde{N}}$ and the state variables are measured, resulting in the dataset
@@
Due to measurement error, the time series data might be contaminated by some noise $\boldsymbol{\epsilon}_{x}$, such that
@@
For simplicity, here we assume that the noise ${\boldsymbol{\epsilon}}_{x}\in\mathbb{R}^{\tilde{N}\times D}$ consists of $\tilde{N}\times D$ independent and identically distributed random variables with Gaussian distributions, with zero means and variance $\boldsymbol{\sigma}_{x}^{2}\in\mathbb{R}^{\tilde{N}\times D}$.
The goal of SINDy is to discover $f$ from the observed time series data $\mathbf{X}$. SINDy postulates that $f$ can be written as a linear combination of a library of candidate functions. For example, a commonly used library is the polynomial library
@@
in which we assume the library consists of $M$ columns of possible candidates. Here, the product and power operations $\mathbf{x}^{n}\mathbf{y}^{m}=[x_{i}^{n}y_{i}^{m}]$ are applied elementwise.
If $f$ can be described by a few terms in the library of candidate functions, then the parameter vector $\mathbf{w}$ that fulfils
@@
will be sparse. The aim of SINDy is to recover this sparse parameter vector $\mathbf{w}$ through regression. If $f$ cannot be fully described by a linear composition of library candidate functions, but can be approximated by it, then SINDy finds a parsimonious approximation to $f$ through sparse regression.
However, in practice, the time derivative is usually not measured as part of the dataset. To recover the time derivative from the time-series data $\mathbf{X}$, we assume there exist some linear operators $L_{I}$ and $L_{\partial t}$ such that
$L_{I}\dot{\mathbf{X}}_{i}\approx L_{\partial t}\mathbf{X},\quad\mbox{and therefore}\quad L_{\partial t}\mathbf{X}\approx L_{I}f(\mathbf{X})$ (5)
in the absence of noise, i.e. when $\boldsymbol{\epsilon}_{x}=0$. Note that $L_{I}$ and $L_{\partial t}$ have dimensions $\hat{N}\times N$. Further discussion on these operators for recovering time derivatives will follow in §2.3.
Plugging (3) and (5) into (4), we rewrite the regression as
@@
Here, we see that the noise is transformed by the operations on $\boldsymbol{\epsilon}_{x}$, giving rise to the total noise $\boldsymbol{\epsilon}=\boldsymbol{\epsilon}_{\Theta}+\boldsymbol{\epsilon}_{\partial t}$, which consists of $\boldsymbol{\epsilon}_{\partial t}$, the noise arising from the time derivative, and $\boldsymbol{\epsilon}_{\Theta}$, the noise arising from the nonlinear library. Since $L_{\partial t}$ is a linear operator on $\mathbf{X}$, assuming $L_{\partial t}^{2}$ represents element-wise squaring of the matrix representing $L_{\partial t}$, the variance of $\boldsymbol{\epsilon}_{\partial t}$ is
$\boldsymbol{\sigma}_{\partial t}^{2}=L_{\partial t}^{2}\boldsymbol{\sigma}_{x}^{2}$ (7)
and $\boldsymbol{\epsilon}_{\partial t}\sim\mathcal{N}(0,\boldsymbol{\sigma}_{\partial t}^{2})$ remains Gaussian. However, $\boldsymbol{\Theta}(\mathbf{X})$ consists of nonlinear functions of $\mathbf{X}$, so the resulting distribution of $\boldsymbol{\epsilon}_{\Theta}$ is non-Gaussian and probably not tractable analytically. Nonetheless, one can trace the variance of each nonlinear function through the algebra of random variables, giving rise to a library of variance,
@@
where each column is associated with each column in the library of candidate functions. Here, we further approximate that the resulting distribution from each column (function) in $\boldsymbol{\Theta}(\mathbf{X})$ are mutually independent, and independent of $\dot{\mathbf{X}}$. Under these assumptions, the variance of the total noise $\boldsymbol{\epsilon}$ can be written as
@@
where $\mathbf{w}^{2}=[w_{ij}^{2}]$ is the elementwise square of $\mathbf{w}$ and $L_{I}^{2}$ the elementwise square of $L_{I}$.
Therefore, according to (9), the noise variance $\boldsymbol{\sigma}^{2}$ depends on the parameter $\mathbf{w}$. This dependency was omitted in previous work [28, 27, 4], but this work will address this dependency in the regression step. It is also worth reiterating that even though we have used the algebra of random variables to keep track of the noise variance, the final distribution of $\boldsymbol{\epsilon}^{2}$ is expected to be non-Gaussian. Nonetheless, we shall later approximate it as a Gaussian distribution in order to preserve analytical tractability.
For simpler notation, we define the dependent variable in (6) as
$\mathbf{t}=L_{\partial t}\mathbf{X}\in\mathbb{R}^{N\times D}$ (10)
and the dictionary, which is the library of functions of the independent variable $\mathbf{X}$, as
@@
Substituting the redefined variables $\mathbf{t}$ and $\mathbf{D}$ into (6), the sparse regression problem becomes
$\mathbf{t}=\mathbf{D}\mathbf{w}+\boldsymbol{\epsilon}$ (12)
in which $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{\beta}^{-1})$ is approximated to be a Gaussian noise with zero mean and elementwise variance $\boldsymbol{\beta}^{-1}=\boldsymbol{\sigma}^{2}$. Here, we use the inverse noise variance $\boldsymbol{\beta}^{-1}$ instead of $\boldsymbol{\sigma}^{2}$ to facilitate easier notation in the next section.

### v2

@@ 1 Introduction
Another prevalent problem with existing SINDy-like algorithms is their ability to handle scarce and noisy data. While SINDy-based approaches are effective at identifying equations from synthetic data with little noise, real-world measurements are typically scarce and prone to significant measurement error. Large noise in the data presents several challenges to SINDy. First, measurement noise corrupts the precise recovery of time derivatives because conventional finite difference approximations can amplify noise by the order of the sample rate. This problem can be mitigated by using a weak formulation [13, 14]. However, this approach may also introduce bias to the data, as we will illustrate in §2.3. Second, excessive noise in the data can impede STLS’s ability to select the most appropriate model. STLS’ modest performance in noisy and scarce data is partly due to the fact that it does not account for the uncertainty in the data, even though this uncertainty can be a crucial factor to consider when choosing a model. Furthermore, the performance of STLS deteriorates when the number of data points is lower than a certain threshold. As we will demonstrate later in §4.1, by adopting a Bayesian viewpoint, one can better find the reason for such deterioration and the remedy against it, effectively addressing the constraints posed by scarce and noisy data.
@@
Therefore, the goal of this study is to demonstrate how a Bayesian perspective can encompass all preceding techniques and is best suited for extending the current SINDy techniques to function in challenging data environments where data are noisy and scarce. As we shall demonstrate later in §4.1, by grounding SINDy in the Bayesian framework, we can also more rigorously justify how and why SINDy works well in the high data low noise limit.
@@ 1.1 Related work and contribution of this work
Sparse Bayesian learning (or regression) is a mature and well-established method in the statistical literature [19, 22, 23]. The dictionary learning technique typically employs priors that promote sparsity. The current ‘gold standard’ sparsity-inducing priors are the spike-and-slab-typed priors, but earlier work has also investigated the use of hierarchical priors. For example, when one combines a Gaussian prior with variance $\alpha^{-1}$ with a maximum-entropy hyperprior on $\alpha$ (i.e. a scale uniform distribution), the result is a sparsity-inducing student-t true prior [24]. A notable early study that utilises this hierarchical prior to promote sparsity is the RVM/SparseBayes package developed by Tippings & Faul [25]. The Relevance Vector Machine (RVM) is a type-II-maximum-likelihood algorithm that was initially created as a probabilistic alternative to the Support Vector Machine (SVM). It fits the data to a dictionary made up of kernels similar to those in SVM, it does so by maximising evidence rather than likelihood, which favours sparsity. This method exploits the fact that the conditional hyperprior given the data can be approximated as a Dirac Delta, while the rest of the probability distributions can be approximated as Gaussians, which allows for analytical tractability. The resulting approach has been demonstrated to be both computationally efficient and robust, not only for classification tasks but also for addressing general sparse regression problems. Due to its Bayesian foundation, it is also resilient to noisy and scarce datasets. Therefore, it is especially appropriate for the objective of this study.
@@
We note that this study is not the first to apply sparse Bayesian regression in the context of equation discovery or system identification [26, 27, 28], nor the first to integrate sparse Bayesian regression with SINDy [29, 4]. For instance, Zhang & Lin [29] incorporated an additional thresholding procedure into the SparseBayes method and applied the technique to extract different ordinary differential equations (ODEs) and partial differential equations (PDEs) from numerical data. They conducted a comparative analysis of their threshold sparse Bayesian regression algorithm with STLS and LASSO, and demonstrated Bayesian regression’s superior performance in the presence of noisy data. Recent work by Niven et al. [28] has also attempted to use Bayesian inference for model selection and uncertainty quantification in system identification. However, the synthetic dataset they considered only introduced uncertainty in the time derivatives and not in the state variables themselves, which can oversimplify the challenge of the system identification problem, as we shall demonstrate later in (8). Fuentes et al. [4] substituted STLS with SparseBayes directly as the sparse regression algorithm and employed the technique for system identification. In particular, apart from synthetic data, they have successfully applied this approach to two experimental benchmarks. Nevertheless, the dataset they used has a high sampling rate (i.e. a high data limit). When finite difference was applied, the noise was significantly amplified in their time derivative, resulting in excessively large uncertainty in their learning result despite successfully recovering the desired equations.
@@
While the above work uses variants of SparseBayes, which do not require expensive computation, they rely on Gaussian approximations of the likelihood and prior distribution in order to maintain the analytical tractability of the distributions. In contrast, subsequent work [30, 31, 32, 33] use spike-and-slab-typed priors, which, despite the more accurate recovery of the correct model, necessitates costly sampling techniques, such as MCMC [30, 32] or Approximated Bayesian Computation [31] to capture the distribution of the estimated parameter more accurately. The benefit of using these sampling-based techniques lies in their capacity to precisely capture non-Gaussian distributions in the parameter and their ability to integrate sparsity-promoting spike-and-slab-typed prior. However, they come at the expense of expensive computations. A cheaper alternative is to make use of ensembling technique [12] to estimate uncertainty in the model. Recent work [34] has demonstrated the convergence of the bootstrapping-based STLS algorithm to the equivalent Bayesian statistics. Nonetheless, these algorithms remain an order of magnitude more expensive than STLS and the Gaussian approximations used by SparseBayes.
@@
Given that our goal is to seek a fast and computationally inexpensive learning method, this study will follow the earlier work more closely [29, 4] and shall adopt strategies similar to the SparseBayes method. Despite its similarity with the earlier work, we note a few key differences. First, our work highlights the key strength of the Bayesian framework, namely, its superior performance in scarce and noisy data. This is particularly in contrast to [4], whose dataset is in the high data limit. Second, our work compares the performance of the weak formulation and finite difference in obtaining derivatives from scarce and noisy data and highlights the pros and cons of each method, which is in contrast to the finite difference used in earlier work [29, 4]. Lastly, we shall present a new sparse Bayesian regression algorithm, which, despite heavy inspiration from SparseBayes, gives better performance due to its ability to keep track of how noise variance propagates throughout the calculation.
@@ 1.2 Structure of the paper
The rest of the paper is structured as follows. In §2, we will provide a concise overview of the SINDy framework. Then, we will introduce our Bayesian approach and estimation used to approximate how noise propagate through the collection of potential functions and temporal derivatives. The limitation of these approximations will be discussed later in §4.2. We will also briefly compare our Bayesian algorithm with previous studies, and highlight the potential bias introduced by the weak formulations when the sampling rate is low, which further reinforces our case for a noise-robust algorithm (§2.2). In §3, we will showcase the efficacy of the suggested framework by presenting examples such as the Van Der Pol system, the nonlinear oscillator, the Lorenz system, and predator-prey dynamics. The predator-prey dynamics are based on real-life data, highlighting the practicality of the framework in real-world applications. Lastly, we will analyse the reasons why the Bayesian approach is more resilient in scarce and noisy data compared to STLS (§4.1), but not as resilient as traditional Bayesian techniques due to the trade-offs made to speed up computation (§4.2). We will also discuss how the improved data efficiency and the probability framework create new potential innovations in active learning (§4.3).
@@ 2.1 SINDy and tracing noise propagation
Suppose we have a system of interest with state variables $\hat{\mathbf{x}}$ with dimension $D$, defined as $\hat{\mathbf{x}}(t):=[\hat{x}_{1}(t),\;\hat{x}_{2}(t)\;...\;\hat{x}_{D}(t)]\in\mathbb{R}^{1\times D}$, where $t$ is time. The dynamics of $\hat{\mathbf{x}}(t)$ is governed by the dynamical equation
@@
with an unknown function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. Suppose the system was observed $\tilde{N}$ times at $t=t_{1},\;t_{2},\;...t_{\tilde{N}}$ and the state variables are measured, resulting in the dataset
@@
Due to measurement error, the data might be contaminated by noise $\boldsymbol{\epsilon}_{x}$, such that the observed variables are:
@@
Here, we assume that the noise ${\boldsymbol{\epsilon}}_{x}\in\mathbb{R}^{\tilde{N}\times D}$ consists of $\tilde{N}\times D$ independent random variables with Gaussian distributions, with zero means and variance $\boldsymbol{\sigma}_{x}^{2}\in\mathbb{R}^{\tilde{N}\times D}$. This simplifiying assumption is introduced mostly for pedagogical simplicity. More discussion on potential extension to the noise model beyond independent Gaussian noise will be give in §4.2.
The goal of SINDy is to discover $f$ from the observed time series data $\mathbf{X}$. SINDy postulates that $f$ can be written as a linear combination of a library of candidate functions. For example, a commonly used library is the polynomial library of the state variables:
@@
in which we assume the library consists of $M$ columns of possible candidates. Here, the product and power operations $\mathbf{x}^{n}\mathbf{y}^{m}=[x_{i}^{n}y_{i}^{m}]$ are applied elementwise. If $f$ can be described by a few terms in the library of candidate functions, then the parameter vector $\mathbf{w}$ that fulfils
@@
will be sparse. The aim of SINDy is to recover this sparse parameter vector $\mathbf{w}$ through regression. If, however, $f$ cannot be fully described by a linear composition of library candidate functions, but can be approximated by it, then SINDy finds a parsimonious approximation to $f$ through sparse regression.
In practice, the time derivative is usually not measured as part of the dataset. To recover the time derivative from the time-series data $\mathbf{X}$, we assume there exist some linear operators $L_{\partial t}$ and $L_{I}$ that approximates the time derivative and the interpolation at $N$ time points, respectively, such that
$L_{I}\dot{\mathbf{X}}\approx L_{\partial t}\mathbf{X},\quad\mbox{and therefore}\quad L_{\partial t}\mathbf{X}\approx L_{I}f(\mathbf{X})$ (5)
in the absence of noise, i.e. when $\boldsymbol{\epsilon}_{x}=0$. Therefore, $L_{I}$ and $L_{\partial t}$ are matrices with dimensions $N\times\tilde{N}$. Further discussion on these operators for recovering time derivatives will follow in §2.3.
Substituting (3) and (5) into (4), we rewrite the regression as
@@
The total noise $\boldsymbol{\epsilon}$ originates from the propagation of $\boldsymbol{\epsilon}_{x}$ through the transformations in (6). It is probably analytically intractable to calculate the full distribution of $\boldsymbol{\epsilon}$, which motivates the use of sampling methods like MCMC. Nevertheless, the variance of $\boldsymbol{\epsilon}$ can be estimated by tracing how the noise variance propagates through (6), which can then be used to guide model selection under the Bayesian framework. Although this variance approximation is crude, it allows for inexpensive computation, which is particularly crucial in real-time applications. In other words, this study foregoes a complete and accurate calculation of the noise distribution in favour of a rapid approximation that enables real-time model selection or discovery. Further analysis of the constraints of this approximation will follow in §4.2.
Here, we outline our method for estimating the variance $\boldsymbol{\sigma}^{2}$ of the total noise $\boldsymbol{\epsilon}$. Firstly, we assume that the statistics of noise arising from each nonlinear function (each column) in $\boldsymbol{\Theta}(\mathbf{X})$ and time derivative $L_{\partial t}{\mathbf{X}}$ are independent. Secondly, we assume that the noise at each time point in (6) is independent of each others, i.e. we approximate the covariance matrix by its diagonal. These approximations circumvent the calculation of correlations between nonlinear terms, which is combinatorially expensive, and between time points, which necessitates the storage and inversion of a large covariance matrix. These are ad hoc approximations aimed at reducing the computational cost, and their theoretical justification may require further investigation. However, we can empirically show their effectiveness in the examples later (e.g. Figure 7).
Because each term is assumed to be independent, we can break down the computation of the noise variance of the total noise $\boldsymbol{\epsilon}$ into contributions from each term in (6). The contribution from the time derivative $\boldsymbol{\sigma}_{\partial t}$ is given by $\boldsymbol{\sigma}^{2}_{\partial t}=L_{\partial t}^{2}\boldsymbol{\sigma}_{x}^{2}$, assuming $L_{\partial t}^{2}$ is the elementwise square of the matrix representing the linear operator $L_{\partial t}$. (Note that the full covariance matrix should be $L_{\partial t}\mbox{diag}(\boldsymbol{\sigma}_{x}^{2})L_{\partial t}^{T}$ for each column of $\boldsymbol{\sigma}_{x}^{2}$. However, since the variance in the diagonal dominates over the off-diagonal terms, we assume temporal independence and neglect the off-diagonal contribution to reduce computational cost.) The contribution from the library can be estimated by a Taylor expansion of statistical moments [35]. Since nonlinear functions are applied elementwise, we can construct a corresponding library of variance,
@@
where each element is the variance of the corresponding element in $\boldsymbol{\Theta}(\mathbf{X})$. The exact formula used to calculate the variance of each term in the library is given in Appendix A. By the above approximations, the variance $\boldsymbol{\sigma}^{2}$ of the total noise $\boldsymbol{\epsilon}$ can be written as
@@
where $\mathbf{w}^{2}=[w_{ij}^{2}]$ is the elementwise square of $\mathbf{w}$ and $L_{I}^{2}$ the elementwise square of $L_{I}$. Notice that the noise variance $\boldsymbol{\sigma}^{2}$ depends on the parameter $\mathbf{w}$. This dependency was omitted in previous work [29, 28, 4], but this study will address this dependency in the regression step.
For simpler notation, we define the left-hand side of the regression in (6) as
$\mathbf{y}=L_{\partial t}\mathbf{X}\in\mathbb{R}^{N\times D}$ (9)
and the dictionary, which is the library of functions of the measured state variables $\mathbf{X}$, as
@@
Substituting the redefined variables $\mathbf{y}$ and $\mathbf{D}$ into (6), the sparse regression problem becomes
$\mathbf{y}=\mathbf{D}\mathbf{w}+\boldsymbol{\epsilon}.$ (11)
As a final approximation, we approximate $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{\beta}^{-1})$ as a Gaussian noise with zero mean and elementwise variance $\boldsymbol{\beta}^{-1}=\boldsymbol{\sigma}^{2}$. Here, we use the inverse noise variance $\boldsymbol{\beta}^{-1}$ instead of $\boldsymbol{\sigma}^{2}$ to facilitate easier notation in the next section. We assume this so that we can employ the following fast Bayesian regression algorithm. The validity of this approximation will be discussed in §4.2.